OpenStudy (anonymous):

What is the vertex of the parabola defined by the equation (x − 2)2 = -12(y − 2)?

OpenStudy (anonymous):

do you have choices?

OpenStudy (anonymous):

(-12, 2) (2, 2) (6, 2) (2, -2

OpenStudy (anonymous):

its b because it shows me on the graph i made

OpenStudy (anonymous):

thanks can u help me with another

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A parabola is defined by the equation x = 5y2. In which direction will the parabola open?

OpenStudy (cjga20049ers):

To find the x-coordinate of the vertex, you use the formula -b/2a. (Remember that the standard form of a quadratic is y = ax2 + bx + c.) So in this case b = 8 and a = 2. -b/2a = -8/(2·2) = -8/4 = -2 Now to find the y-coordinate, simply plug the x-value into the quadratic equation. y = 2x2 + 8x + 9 y = 2(-2)2 + 8(-2) + 9 y = 8 - 16 + 9 y = 1 So the vertex is at the point (-2, 1) choice D 8/11/2013 | CJ